1. You have the chance to participate in a project that produces the following cash flows:...

1.

You have the chance to participate in a project that produces the following cash flows:

Cash Flows ($)
C0 C1 C2
4,600 4,400 –10,800

a. The internal rate of return is 12.69%. If the opportunity cost of capital is 12%, what is the NPV of the project? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 2 decimal places.)

NPV            $ __________.

2.

Consider the following projects:

Cash Flows ($)
Project C0 C1 C2 C3 C4 C5
A −2,500 2,500 0 0 0 0
B −5,000 2,500 2,500 5,500 2,500 2,500
C −6,250 2,500 2,500 0 2,500 2,500

a. If the opportunity cost of capital is 9%, which project(s) have a positive NPV?

Positive NPV project(s)

Project A
Project B
Project C
Projects A and B
Projects A and C
Projects B and C
Projects A, B, and C
No project

b. Calculate the payback period for each project: (Round your answers to 2 decimal places. If a project never pays back, enter "0".)

  Project A year(s)
  Project B year(s)
  Project C year(s)

c. Which project(s) would a firm using the payback rule accept if the cutoff period were three years?

Project(s) accepted                  (Click to select)Project AProject BProject CProjects A and BProjects A and CProjects B and CProjects A, B, and CNo project

d. Calculate the discounted payback for each project. (Do not round intermediate calculations. Round your answers to 2 decimal places. If a project never pays back, enter "0".)

Project A year(s)
Project B year(s)
Project C year(s)

e. Which project(s) would a firm using the discounted payback rule accept if the cutoff period were three years?

Project(s) accepted                (Click to select)Project AProject BProject CProjects A and BProjects A and CProjects B and CProjects A, B, and CNo project

3.

The Titanic Shipbuilding Company has a noncancelable contract to build a small cargo vessel. Construction involves a cash outlay of $253,000 at the end of each of the next two years. At the end of the third year the company will receive payment of $645,000. Assume the IRR of this option exceeds the cost of capital.

The company can speed up construction by working an extra shift. In this case, there will be a cash outlay of $555,000 at the end of the first year, followed by a cash payment of $645,000 at the end of the second year. Use the IRR rule to show the (approximate) range of opportunity costs of capital at which the company should work the extra shift. (Enter your answers as a percent rounded to 2 decimal places. Enter the smallest percent first.)

The company should work the extra shift if the cost of capital is between ______% and ________%

4.

Consider the following projects:

Cash Flows ($)
Project C0 C1
D –11,300 22,600
E –21,300 37,275

Assume that the projects are mutually exclusive and that the opportunity cost of capital is 9%.

a. Calculate the profitability index for each project. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Project Profitability Index
D
E

b-1. Calculate the profitability-index using the incremental cash flows. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Profitability-index _______

b-2. Which project should you choose?

Project D
Project E
Answer #1

1.

Let the cash flow for period n be Cn

C0 = $4600
C1 = $4400
C2 = -$10800

Cost of Capital = r = 12%

Hence, NPV = ΣCn/(1+r)n = C0 + C1/(1+r) + C2/(1+r)2

= 4600 + 4400/1.12 - 10800/1.122 = -$81.12